Shark and Co. purchases products from 3 companies. The expected number of flaws in an order is 4 if the items come from company A, 2.5 if the order comes from company B and 3 if it comes from company C. What is the expected number of flaws in the order, if the order is from company A with probability 0.5, from company B with probability 0.3 and from company C with probability 0.2?
Answer #1

Solution:

Probability distribution table can be constructed as

Company No. of Flaws Probability
A 4 0.5
B 2.5 0.3
C 3 0.2


Expected No. of flaws = Summation(Xi*P(Xi)) = (4*0.5) + (2.5*0.3) + (3*0.2) = 2 + 0.75 + 0.6 = 3.35
Expected No. of Flaws = 3.35